OpenStudy (anonymous):

solve this system of linear equations 9x-19y=74 3x+3y=6

OpenStudy (johnweldon1993):

You could use the elimination method...multiply that bottom equation by 3... 9x - 19y = 74 3(3x + 3y = 6) Now you'll have 9x - 19y = 74 9x + 9y = 18 Now subtract your 2 equations.... 9x - 19y = 74 -(9x + 9y = 18) Now you see that 9x - 9x = 0...so you only have the 'y' variable to solve for... 9x - 9x = 0 -19y - 9y = -28y 74 - 18 = 56 So now we have -28y = 56 Divide both sides by -28....what does 'y' equal??

OpenStudy (anonymous):

2

OpenStudy (johnweldon1993):

Almost....is it 2 or is it -2 ?? 56 / -28 = ...?

OpenStudy (anonymous):

-2

OpenStudy (anonymous):

is x 0

OpenStudy (johnweldon1993):

Right...so now that we know what 'y' equals...take that -2...and substitute it in for the 'y' value in either of your original equations. lets do... 9x - 19y = 74 now replace 'y' with -2 9x - 19(-2) = 74 Simplify a bit... 9x + 38 = 74 Subtract 38 from both sides of the equation 9x = 36 Divide both sides by 9....x = ?

OpenStudy (johnweldon1993):

So yes you'll see that x = 4....and we found that y = -2 Now lets test it.... 9x-19y=74 9(4) - 19(-2) = 74 36 + 38 = 74 74 = 74 so this one is correct....now lets check the other... 3x+3y=6 3(4) + 3(-2) = 6 12 - 6 = 6 6 = 6 They are both correct...so you have the correct answers